It is necessary to have a 40% antifreeze solution in the radiator of the certain car. The radiator now has 70liters of 20% solution. How many liters of this should be drained and replaced with 100% Antifreeze to get the desired strength?
Answer by Paul(988)

-100x+20(70)=40(-x+20)
-100x+1400=-40x+800
-60x = -600
x=10

Hence about 10 liters should be drained.
Paul.